You are planning to make monthly deposits of $150 into a retirement account that pays 14 percent interest compounded monthly. If your first deposit will be made one month from now, your retirement account will be worth $___in 15 years
Answer #1

Rate = 14% / 12 = 1.16667%

Number of periods = 15 * 12 = 180

Future value = Annuity * [(1 + r)n - 1] / r

Future value = 150 * [(1 + 0.0116667)180 - 1] / 0.0116667

Future value = 150 * 605.788643

Future value = $90,867.94